BRICS Lecture Series, 1995–98

March 4, 2002

This document is also available as PostScript, DVI, Text.

References

[LS-01-1] Abstract, PostScript, PDF, DVI.

Zsolt Tuza. Unsolved Combinatorial Problems, Part I. May 2001. viii+30 pp.

[LS-98-4] Abstract, PostScript, PDF, DVI.

Paola Quaglia. The π-Calculus: Notes on Labelled Semantics. December 1998. viii+16 pp.

[LS-98-3] Abstract.

Olivier Danvy. Type-Directed Partial Evaluation. December 1998. Extended version of lecture notes to appear in.

[LS-98-2] Abstract, PostScript, PDF, DVI.

Carsten Butz. Regular Categories and Regular Logic. October 1998.

[LS-98-1] Abstract, PostScript, PDF, DVI.

Ulrich Kohlenbach. Proof Interpretations. June 1998.

[LS-97-1] Abstract, PostScript, PDF, DVI.

Jan Chomicki and David Toman. Temporal Logic in Information Systems. November 1997. viii+42 pp. Full version to appear in: Logics for Database and Information Systems, Chomicki and Saake (eds.), Kluwer Academic Publishers, 1998.

[LS-96-6] Abstract, PostScript, PDF, DVI.

Torben Braner. Introduction to Linear Logic. December 1996. iiiv+55 pp.

[LS-96-5] Abstract, PostScript, PDF.

Devdatt P. Dubhashi. What Can’t You Do With LP? December 1996. viii+23 pp.

[LS-96-4] Abstract, PostScript, DVI.

Sven Skyum. A Non-Linear Lower Bound for Monotone Circuit Size. December 1996. viii+14 pp.

[LS-96-3] Abstract, PostScript.

Kristoffer H. Rose. Explicit Substitution – Tutorial & Survey. September 1996. v+150 pp.

[LS-96-2] Abstract, PostScript, DVI.

Susanne Albers. Competitive Online Algorithms. September 1996. iix+57 pp.

[LS-96-1] Abstract, PostScript, PDF. Lars Arge. External-Memory Algorithms with Applications in Geographic Information Systems. September 1996. iix+53 pp.

[LS-95-5] Abstract, PostScript, DVI. Devdatt P. Dubhashi. Complexity of Logical Theories. September 1995. x+46 pp.

[LS-95-4] Abstract, PostScript.

Dany Breslauer and Devdatt P. Dubhashi. Combinatorics for Computer Scientists. August 1995. viii+184 pp.

[LS-95-3] Abstract, PostScript, DVI. Michael I. Schwartzbach. Polymorphic Type Inference. June 1995. viii+24 pp.

[LS-95-2] Abstract, PostScript, DVI. Sven Skyum. Introduction to Parallel Algorithms. June 1995. viii+17 pp. Second Edition.

[LS-95-1] Abstract, PostScript, DVI. Jaap van Oosten. Basic Category Theory. January 1995. vi+75 pp.